# KCL-PH-TH/2026-03

### v2

@@ V Stringy-Running-Vacuum-Model Inflation
where we expanded to linear order in $\lambda$. Eq. (142) can be solved to yield at a time $t_{\rm begin}<t<t_{\rm end}$ during inflation, where $t_{\rm begin}$ ($t_{\rm end}$) denotes the onset (end) of the inflationary era:
@@
given that $3.04\,\frac{\lambda}{\kappa^{2}}\ll\kappa^{-2}\,H^{-2}\sim 10^{10}$, due to (IV), (120).
This provides a justification for the (approximate) constancy of $\mathcal{K}^{0}$ employed in our argumentation above. The initial value then of $\mathcal{K}^{0}(t_{\rm begin})$ at the onset of the RVM inflation in (143) is determined from (122) by using the parametrization (123) on the left hand side, along with the dynamical-system-analysis estimate of the order of the parameter $\varepsilon=\mathcal{O}(10^{-2})$ [8].
We thus observe that the higher-order $\lambda$-dependent corrections in the effective action (102) do not affect at all the lower bound of $\mathcal{N}_{I}$, as compared to the $\lambda=0$ case, due to the extreme weakness of those terms. Eq. (V) is compatible with the corresponding result in [7], based on a dynamical system analysis, where one finds:
@@
which ensures continuity of the value of the Chern-Simons anomaly condensate during the transition from the stiff to inflationary eras ($\mathcal{N}_{\rm stiff}$ denotes the assumed numerical density of sources during the stiff era that precedes the inflationary epoch in the StRVM [2, 4, 8]). The equality in (V) is satisfied in order of magnitude by $\mathcal{N}_{\rm stiff}=\mathcal{O}(10^{-1})$. So the conclusions of the previous approach of [2, 7] remain qualitatively and quantitatively correct, upon the validity of (120), which is supported by the Planck-data phenomenology of inflation [42].
A final remark we would like to make concerns the consistency of the condensate inflation with classical field theory considerations. Indeed, despite the fact that the condensate is the result of quantum GW primordial excitations, which condense, nonetheless its value remains consistent with the Euler-Lagrange (classical) equation of motion for the axion-like field $b$, after the formation of the condensate, in a FLRW background:
@@
where in the last line we ignored the subleading $\lambda$-dependent contributions. To check the validity of this equation, during the condensate-induced inflation, we rewrite it, for (approximately) constant, i.e. $\dot{b}=0$, as:
@@
Using (120), we obtain from (147):
@@
which is remarkably consistent with the bound (V), if one takes into account the theoretical uncertainties involved, due to the dominant rôle of GW modes near the UV cutoff $M_{s}$ in the formation of the gCS anomaly condensate (130).
The above result indicates that higher-derivative and higher-than-quadratic curvature terms in the string effective action, which were naively expected to play an important rôle for modes with momenta near the UV cutoff, will not affect (in order of magnitude) the analysis of [2, 4, 7] for a primordial GW-condensate-induced inflation, which was based on lowest non-trivial order $\mathcal{O}(\alpha^{\prime})$ string effective actions with gravitational anomalies. The consistency of the condensate approach from many, rather distinct, viewpoints, ranging from dynamical-system analysis, to weak quantum-gravity effective field theory, offers a non-trivial support to this.
@@ VI Conclusions
In this work, we have revisited the String-inspired Running Vacuum Model (StRVM) of Cosmology, proposed in [2, 4, 5, 7, 8], by considering the most general low-energy effective gravitational action quartic in spacetime derivatives of graviton and antisymmetric-tensor fields (assuming a stabilised dilaton to a constant value, which sets the scale for the string coupling, and hence the relevant phenomenology). In (3+1)-dimensional spacetimes we have shown that it is possible to impose simultaneously unitarity (i.e. a Gauss-Bonnet-quadratic curvature scheme, which ensures absence of graviton ghosts), and torsional interpretation of the field strength of the antisymmetric tensor field, which behaves as a totally antisymmetric component of spacetime torsion.
By studying carefully the potential field redefinitions leading to the above features, we have managed to arrive at an appropriate basis of structures, which contains only one extra term of fourth-order in derivatives, as compared to the effective action of the StRVM, of the form
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{\lambda}^{\rm extra}=\lambda\,\int d^{4}x\,\sqrt{-g}\,\partial_{\mu}b\,\partial_{\nu}b\,R^{\mu\nu}\,,\qquad\lambda\propto\alpha^{\prime}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,$,
where $R_{\mu\nu}$ denotes the torsion-free Ricci tensor of the spacetime geometry, and $b$ is the massless axion-like field. The latter is dual (in the (3+1)-dimensional spacetime after string compactification) to the field strength (and totally antisymmetric torsion) $\mathcal{H}_{\mu\nu\rho}$ in the sense of (90), when order $\alpha^{\prime}$ terms in the effective action are taken into account. Nonetheless, for Einstein spaces (as is the inflationary spacetime of interest here) this modified duality relation maintains the essential structure of (1), but now includes a global proportionality scalar factor, dependent on the scalar spacetime curvature (cf. (93)),
@@
We have explained in the main text the way the above $\lambda$-dependent term $\mathcal{S}_{\lambda}^{\rm extra}$ arises, after path-integrating the torsion $\mathcal{H}_{\mu\nu\rho}$ field in the partition function with respect to the generalised four-derivative effective action (52). The axion field is introduced in the path integral as a Bianchi-identity constraint-implementing Lagrange multiplier field, which is originally non-dynamical. The dynamics of $b(x)$ arises after the (Gaussian) path-integration over the $\mathcal{H}$-torsion.
The parameter $\lambda\propto\alpha^{\prime}=M_{s}^{-2}$, and this implies that the contribution of the parity-even term in the parity-odd gravitational anomaly condensate is negligible, for the values of string scale $M_{s}$ required for the correct phenomenology of the StRVM inflation [7]. This leads to the conclusion that the analysis of the StRVM, which ignored such $\lambda$-dependent contributions was complete, in view of the extreme suppression of these terms.
We have also discussed a remarkable consistency of the anomaly-condensate-induced inflation in the context of the StRVM, which stems perhaps from the exactness of the anomaly term per se. Indeed, we have shown how the value of the gravitational anomaly condensate, as estimated by a weak-chiral-GW quantum tensor perturbations effective field theory, is consistent with the satisfaction of the classical Euler-Lagrange equations of motion of the axion $b(x)$ field, thereby pointing towards a “classical” nature of the anomaly condensate. As discussed in [2], this is also consistent with a spontaneous violation of Lorentz symmetry due to the formation of a vacuum epxectation value of the temporal component of the topological current corresponding to the Chern-Simons gravitational anomaly term, as a consequence of the GW-induced anomaly condensation.
Acknowledgements
NEM thanks the University of Valencia and its Theoretical Physics Department for a visiting Research Professorship supported by the programme Atracción de Talento INV25-01-15, during which initiation of this work took place. The research of NEM is supported in part by the UK Science and Technology Facilities research Council (STFC) under the research grants ST/X000753/1. He also acknowledges support by the Severo Ochoa project CEX2023-001292-S, and participation in the COST Association Actions CA21136 “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)” and CA23130 “Bridging high and low energies in search of quantum gravity (BridgeQG)”.
@@ Appendix A Decomposition of the generalized (contorted) curvature tensors and useful identities
Let us assume a torsionful connection of the form:
@@
We can then use the above to “decompose” the generalized curvature (torsionful Riemann) tensor into the usual (Levi-Civita connection) Riemann tensor and the totally antisymmetric torsion $\mathcal{H}$ using the relations:
@@
For a totally antisymmetric torsion tensor, the first Bianchi identity for the Levi-Civita Riemann tensor, $R_{\mu[\nu\lambda\rho]}=0$, results in:
@@
The Bianchi identity for the Kalb-Ramond field strength, $\nabla_{[\rho}\mathcal{H}_{\mu\nu\lambda]}=0$, results in:
@@
$\begin{split}\int\nabla_{\rho}\mathcal{H}_{\mu\nu\lambda}\nabla^{\rho}\mathcal{H}^{\mu\nu\lambda}\sqrt{-g}\,d^{4}x&=\int\nabla_{\rho}(\mathcal{H}_{\mu\nu\lambda}\nabla^{\rho}\mathcal{H}^{\mu\nu\lambda}-3\mathcal{H}^{\mu\rho\nu}\nabla^{\lambda}\mathcal{H}_{\mu\nu\lambda})\sqrt{-g}\,d^{4}x\\ &+3\int\nabla_{\lambda}\mathcal{H}^{\mu\nu\lambda}\nabla^{\rho}\mathcal{H}_{\mu\nu\rho}\sqrt{-g}\,d^{4}x-3\int\mathcal{H}_{\mu}{}^{\lambda\rho}\mathcal{H}_{\nu\lambda\rho}R^{\mu\nu}\sqrt{-g}\,d^{4}x\\ &+3\int\mathcal{H}_{\kappa}{}^{\mu\nu}\mathcal{H}^{\kappa\lambda\rho}R_{\mu\nu\lambda\rho}\sqrt{-g}\,d^{4}x\end{split}$ (159)
Appendix B Field redefinitions directly in D=4 field theory effective actions, beyond string theory
In this Appendix, for completeness, we shall give the result of the field redefinitions (III), (III) directly to a $D=4$ field theory action, which has formally the same form as the $\mathcal{O}(\alpha^{\prime})$ string action (8) ((9), (10)) but here it represents a four-dimensional four-derivative field theory action, independent of strings, for which the standard S-matrix equivalence theorems [33, 34, 35, 36, 37, 38] apply. In such a case the parameter
@@ Appendix B Field redefinitions directly in D=4 field theory effective actions, beyond string theory
may be identified with the inverse of an energy scale below which the effective field theory is valid, i.e. an UV cutoff, not related to the string case. Below we shall outline the relevant differences between the two formalisms, and show that in the context of a cosmological model, like the StRVM, the basic conclusions are not affected in order of magnitude.
@@
and matching with (45), yields
@@
Since in $D=4$ the Gauss-Bonnet quadratic-curvature combination is a total derivative, which is thus ignored, the full effective action reads:
$S=\int\left[\frac{1}{2\kappa^{2}}R-\frac{1}{6}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}+\alpha^{\prime}\left(\frac{1}{36\sqrt{3}}-\frac{1}{16}\right)\left(\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{1}{3}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\right)\right]\sqrt{-g}\,d^{4}x\,.$ (167)
The reader can compare directly (167) with (52), and the difference arises in both, the magnitude and signature of, the coefficient of the $\mathcal{O}(\alpha^{\prime})$ terms.
The resulting effective action (in terms of the axion fields, that arise after the $\mathcal{H}$-torsion path integration) maintains its generic form (102), but now with
@@
where now $\alpha^{\prime}$ is given by (160), related to the UV cutoff scale of the effective theory, which is not the string scale. As this is a gravitational theory, it is natural to identify $\mathcal{M}$ with the effective Planck scale $q$ (B). Due to the positive nature of the parameter $\lambda$ in this case, we observe that the effective Planck energy scale $q^{-1}$ is smaller than the bare one, $\kappa^{-1}$. In fact to maintain unitarity of the theory ($q^{2}>0$) we need to impose the constraint:
@@
which is naturally satisfied if we impose the transplanckian conjecture, according to which no energy scale in our effective field theory should exceed the Planck scale. The opposite happens in section IV, (IV), where, because of the negative nature of the respective parameter $\lambda$, the effective Planck energy scale $q$ (which is always positive for any $\lambda<0$) is larger than the bare $\kappa^{-1}$.
In the context of the D=4 dimensional StRVM, with the improved effective action (102), the value of $0<\lambda=0.093\frac{1}{\mathcal{M}^{2}}$ is such that, with the natural identification $\mathcal{M}=M_{\rm Pl}=q^{-1}$, the $\lambda$-dependent term in (102) still makes subleading contributions to the gCS condensate in the context of the StRVM, for which (123), (120) are valid.
Appendix C Parity-odd quadratic-curvature invariants in 4D
In the context of our study of quadratic generalized curvature invariants, it is worth making a short comment about parity-odd quadratic invariants. These are invariants that involve two generalized curvature tensors and the Levi-Civita tensor. Thus, their existence, number and explicit form are strongly dependent on the number of spacetime dimensions. It is trivial to see, for example, that such invariants do not exist in any odd number of dimensions, since the number of indices in those is odd and a scalar cannot be formed. In even number of dimensions, such invariants exist up until $D=8$, because that is the maximum number of free indices two generalized curvature tensors have. Therefore, the absence of such invariants in $10D$ Heterotic string action should not be a surprise, since they simply do not exist. In $D=4$, the number of parity-odd invariants is four [39]:

### v3

@@ V Stringy Running Vacuum Model Inflation
$\simeq\mathcal{N}_{I}\,(1-2\lambda\kappa^{2}\dot{b}^{2})\,(1-3\lambda H^{2})\,\,\frac{\epsilon^{2}}{64\,\pi^{2}}\,\kappa^{4}\,\mu^{4}\,H^{3}\,$
@@
where we expand to linear order in $\lambda$. Equation (143) can be solved to yield at a time $t_{\rm begin}<t<t_{\rm end}$ during inflation, where $t_{\rm begin}$ ($t_{\rm end}$) denotes the onset (end) of the inflationary era:
@@
given that $3.04\,\frac{\lambda}{\kappa^{2}}\ll\kappa^{-2}\,H^{-2}\sim 10^{10}$ due to (IV) and (121).
This provides a justification for the approximate constancy of $\mathcal{K}^{0}$ employed in our argumentation above. The initial value then of $\mathcal{K}^{0}(t_{\rm begin})$ at the onset of the RVM inflation in (144) is determined from (123) by using the parametrization (124) on the left-hand side, along with the dynamical system analysis estimate of the order of the parameter $\varepsilon=\mathcal{O}(10^{-2})$ [8].
We thus observe that the higher-order $\lambda$-dependent corrections in the effective action (102) do not affect at all the lower bound of $\mathcal{N}_{I}$, as compared to the $\lambda=0$ case, due to the extreme weakness of these terms. Equation (V) is compatible with the corresponding result in [7] based on a dynamical system analysis, where one finds
@@
which ensures continuity of the value of the Chern–Simons anomaly condensate during the transition from the stiff to inflationary eras ($\mathcal{N}_{\rm stiff}$ denotes the assumed numerical density of sources during the stiff era that precedes the inflationary epoch in the StRVM [2, 4, 8]). The equality in (V) is satisfied in order of magnitude by $\mathcal{N}_{\rm stiff}=\mathcal{O}(10^{-1})$. So, the conclusions of the previous approach of [2, 7] remain qualitatively and quantitatively correct, upon the validity of (121), which is supported by the Planck data phenomenology of inflation [44].
A remark we would like to make at this point concerns the consistency of the condensate inflation with classical field theory considerations. Indeed, despite the fact that the condensate is the result of quantum GW primordial excitations, which condense, its value remains nonetheless consistent with the Euler–Lagrange (classical) equation of motion for the axion-like field $b$ after the formation of the condensate in an FLRW background:
@@
where in the last line, we ignored the subleading $\lambda$-dependent contributions. To check the validity of this equation, during the condensate-induced inflation, we rewrite it for an approximate constant, i.e., $\dot{b}=0$, as follows:
@@
Using (121), we obtain the following from (148):
@@
which is remarkably consistent with the bound (V), if one takes into account the theoretical uncertainties involved, due to the dominant role of GW modes near the UV cutoff $M_{s}$ in the formation of the gCS anomaly condensate (131).
The above result indicates that higher-derivative and higher-than-quadratic curvature terms in the string effective action, which were naively expected to play an important role for modes with momenta near the UV cutoff, will not affect (in order of magnitude) the analysis of [2, 4, 7] for a primordial GW condensate-induced inflation, based on lowest-non-trivial-order $\mathcal{O}(\alpha^{\prime})$ string effective actions with gravitational anomalies. The consistency of the condensate approach from many rather distinct viewpoints, ranging from dynamical system analysis to weak quantum gravity effective field theory, offers a non-trivial support to this.
Finally, before concluding this section, we remark that as shown in [8], one can achieve an extremely good agreement in the phenomenology of the gCS condensate inflationary scenario described above with the standard scale invariance violating small slow-roll inflationary parameters, $n_{s},\varepsilon$, and $\eta$, as defined in the standard inflationary phenomenology [44] $\varepsilon_{1}=\frac{1}{2\,\kappa^{2}}\Big(\frac{V^{\prime}}{V}\Big)^{2}$, with the prime denoting a derivative with respect to the field $b$, $\eta=\frac{1}{\kappa^{2}}\,\frac{V^{\prime\prime}}{V}$, and the running spectral index $n_{s}=1-6\,\varepsilon_{1}+2\,\eta$. In addition, the ratio of the scalar to tensor perturbations $r=16\,\varepsilon$. Indeed, by considering instanton effects in the non-Abelian gauge sector that may characterize the model, one obtains an effective, instanton-induced periodic modulation of the axion $b$ potential as follows:
$V(b)_{\rm eff}\ni\Lambda_{1}^{4}\,\cos\Big(2\,\pi^{2}\,\epsilon\,b(x)\Big)\equiv\Lambda_{1}^{4}\,\cos\Big(\frac{b}{f_{b}}\Big)\,$, (150)
where $\Lambda_{1}$ is the instanton energy scale, and $\epsilon$ is defined in (IV). The quantity $f_{b}$ is the axion $b$ coupling, defined through the gauge sector of the Prontryagin anomaly term in the action [24], upon the inclusion of the gauge sector, as follows [8]:
$\mathcal{S}_{\rm anom}\ni\frac{1}{16\pi^{2}\,f_{b}}\int d^{4}x\sqrt{-g}\,b(x)\,{\mathbf{F}}_{\mu\nu}\,\widetilde{{\mathbf{F}}}^{\mu\nu}\,$, (151)
where ${\mathbf{F}}_{\mu\nu}$ is the non-Abelian-gauge-group field strength and $\frac{1}{16\pi^{2}}\int d^{4}x\sqrt{-g}\,{\mathbf{F}}_{\mu\nu}\,\widetilde{{\mathbf{F}}}^{\mu\nu}=n$, $n\in\mathbb{Z}$, the Pontryagin index [24]. By tuning appropriately the scale $\Lambda_{1}$ to certain natural values and assuming $\frac{\kappa}{\sqrt{\alpha^{\prime}}}={\mathcal{O}}(0.1)$, as dictated by the inflationary phenomenology of the model of [7] (i.e., the requirement that the duration of inflation be in the region of 607060-70 e-foldings), we obtain [8], a per-mil agreement on the values of the slow-roll parameters of the model with those measured by the Planck collaboration [44].
@@ VI Conclusions
In this work, we revisited the string-inspired running vacuum model (StRVM) of cosmology proposed in [2, 4, 5, 7, 8] by considering the most general low-energy effective gravitational action quartic in spacetime derivatives of graviton and antisymmetric tensor fields (assuming a stabilized dilaton to a constant value, which sets the scale for the string coupling, and hence the relevant phenomenology). In (3+1)-dimensional spacetimes, we have shown that it is possible to simultaneously impose unitarity (i.e., a Gauss–Bonnet-quadratic curvature scheme, which ensures absence of graviton ghosts) and torsional interpretation of the field strength of the antisymmetric tensor field, which behaves as a totally antisymmetric component of spacetime torsion.
By carefully studying the potential field redefinitions leading to the above features, we managed to arrive at an appropriate basis of structures that contains only one extra term of the fourth-order in derivatives, as compared to the effective action of the StRVM, of the form
$\mathcal{S}_{\lambda}^{\rm extra}=\lambda\,\int d^{4}x\,\sqrt{-g}\,\partial_{\mu}b\,\partial_{\nu}b\,R^{\mu\nu}\,,\qquad\lambda\propto\alpha^{\prime}\,$,
where $R_{\mu\nu}$ denotes the torsion-free Ricci tensor of the spacetime geometry, and $b$ is the massless axion-like field, which, as discussed in Section (IV) (cf. (105)), is expected from the seminal work of [42, 43]. The axion field $b(x)$ is dual (in the (3+1)-dimensional spacetime after string compactification) to the field strength (and totally antisymmetric torsion) $\mathcal{H}_{\mu\nu\rho}$ in the sense of (90) when order $\alpha^{\prime}$ terms in the effective action are taken into account. Nonetheless, for Einstein spaces (as is the inflationary spacetime of interest here), this modified duality relation maintains the essential structure of (1), but now includes a global proportionality scalar factor, dependent on the scalar spacetime curvature (cf. (54), (93)), as follows:
@@
We explained in the main text the way the above $\lambda$-dependent term $\mathcal{S}_{\lambda}^{\rm extra}$ arises after path-integrating the torsion $\mathcal{H}_{\mu\nu\rho}$ field in the partition function with respect to the generalized four-derivative effective action (III). The axion field is introduced in the path integral as a Bianchi identity constraint implementing a Lagrange multiplier field, which is originally non-dynamical. The dynamics of $b(x)$ arise after the Gaussian path integration over the $\mathcal{H}$-torsion.
The parameter $\lambda\propto\alpha^{\prime}=M_{s}^{-2}$, and this implies that the contribution of the parity-even term in the parity-odd gravitational anomaly condensate is negligible for the values of string scale $M_{s}$ required for the correct phenomenology of the StRVM inflation [7]. This leads to the conclusion that the analysis of the StRVM, which ignored such $\lambda$-dependent contributions, is complete in view of the extreme suppression of these terms.
We also discussed a remarkable consistency of the anomaly condensate-induced inflation in the context of the StRVM, which stems perhaps from the exactness of the anomaly term per se. Indeed, we have shown how the value of the gravitational anomaly condensate, as estimated by weak-chiral-GW quantum tensor perturbation effective field theory, is consistent with the satisfaction of the classical Euler–Lagrange equations of motion of the axion $b(x)$ field, thereby pointing towards a “classical” nature of the anomaly condensate. As discussed in [2], this is also consistent with a spontaneous violation of Lorentz symmetry due to the formation of a vacuum epxectation value of the temporal component of the topological current corresponding to the Chern–Simons gravitational anomaly term as a consequence of the GW-induced anomaly condensation.
@@ Appendix A Decomposition of the Generalized (Contorted) Curvature Tensors and Useful Identities
Let us assume a torsionful connection of the following form:
@@
We can then use the above to “decompose” the generalized curvature (torsionful Riemann) tensor into the usual (Levi–Civita connection) Riemann tensor and the totally antisymmetric torsion $\mathcal{H}$ using the following relations:
@@
For a totally antisymmetric torsion tensor, the first Bianchi identity for the Levi–Civita Riemann tensor, $R_{\mu[\nu\lambda\rho]}=0$, results in
@@
The Bianchi identity for the Kalb–Ramond field strength, $\nabla_{[\rho}\mathcal{H}_{\mu\nu\lambda]}=0$, results in
@@
$\begin{split}\int\nabla_{\rho}\mathcal{H}_{\mu\nu\lambda}\nabla^{\rho}\mathcal{H}^{\mu\nu\lambda}\sqrt{-g}\,d^{4}x&=\int\nabla_{\rho}(\mathcal{H}_{\mu\nu\lambda}\nabla^{\rho}\mathcal{H}^{\mu\nu\lambda}-3\mathcal{H}^{\mu\rho\nu}\nabla^{\lambda}\mathcal{H}_{\mu\nu\lambda})\sqrt{-g}\,d^{4}x\\ &+3\int\nabla_{\lambda}\mathcal{H}^{\mu\nu\lambda}\nabla^{\rho}\mathcal{H}_{\mu\nu\rho}\sqrt{-g}\,d^{4}x-3\int\mathcal{H}_{\mu}{}^{\lambda\rho}\mathcal{H}_{\nu\lambda\rho}R^{\mu\nu}\sqrt{-g}\,d^{4}x\\ &+3\int\mathcal{H}_{\kappa}{}^{\mu\nu}\mathcal{H}^{\kappa\lambda\rho}R_{\mu\nu\lambda\rho}\sqrt{-g}\,d^{4}x\,.\end{split}$ (162)
Appendix B Field Redefinitions Directly in D = 4 Field Theory Effective Actions Beyond String Theory
In this Appendix, for completeness, we shall give the result of the field redefinitions (III) and (III) directly to a $D=4$ field theory action, which formally has the same form as the $\mathcal{O}(\alpha^{\prime})$ string action (8) ((9), (10)) but, here, represents a four-dimensional four-derivative field theory action, independent of strings, for which the standard S-matrix equivalence theorems [33, 34, 35, 36, 37, 38] apply. In such a case, the parameter
@@ Appendix B Field Redefinitions Directly in D = 4 Field Theory Effective Actions Beyond String Theory
may be identified with the inverse of an energy scale below which the effective field theory is valid, i.e., an UV cutoff, not related to the string case. Below, we shall outline the relevant differences between the two formalisms and show that in the context of a cosmological model, like the StRVM, the basic conclusions are not affected in order of magnitude.
@@
and matching with (45) yields
@@
Since in $D=4$ the Gauss–Bonnet quadratic curvature combination is a total derivative, which is thus ignored, the full effective action reads
$S=$ $\int\bigg[\frac{1}{2\kappa^{2}}R-\frac{1}{6}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}$
$+\alpha^{\prime}\left(\frac{1}{36\sqrt{3}}-\frac{1}{16}\right)\left(\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{1}{3}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\right)\bigg]\sqrt{-g}\,d^{4}x\,.$ (170)
The reader can compare directly (B) with (III), and the difference arises in both the magnitude and signature of the coefficient of the $\mathcal{O}(\alpha^{\prime})$ terms.
The resulting effective action (in terms of the axion fields that arise after the $\mathcal{H}$-torsion path integration) maintains its generic form (102), but now with
@@
where now $\alpha^{\prime}$ is given by (163), related to the UV cutoff scale of the effective theory, which is not the string scale. As this is a gravitational theory, it is natural to identify $\mathcal{M}$ with the effective Planck scale $q$ (B). Due to the positive nature of the parameter $\lambda$ in this case, we observe that the effective Planck energy scale $q^{-1}$ is smaller than the bare one, $\kappa^{-1}$. In fact, to maintain unitarity of the theory ($q^{2}>0$), we need to impose the following constraint:
@@
This constraint is naturally satisfied if we impose the transplanckian conjecture, according to which no energy scale in our effective field theory should exceed the Planck scale. The opposite happens in Section IV, (IV), where, because of the negative nature of the respective parameter $\lambda$, the effective Planck energy scale $q$ (which is always positive for any $\lambda<0$) is larger than the bare $\kappa^{-1}$.
In the context of the D=4 dimensional StRVM, with the improved effective action (102), the value of $0<\lambda=0.093\frac{1}{\mathcal{M}^{2}}$ is such that with the natural identification $\mathcal{M}=M_{\rm Pl}=q^{-1}$, the $\lambda$-dependent term in (102) still makes subleading contributions to the gCS condensate in the context of the StRVM, for which (124) and (121) are valid.
Appendix C Parity-Odd Quadratic Curvature Invariants in 4D
In the context of our study of quadratic generalized curvature invariants, it is worth making a short comment about parity-odd quadratic invariants. These are invariants that involve two generalized curvature tensors and the Levi–Civita tensor. Thus, their existence, number, and explicit form are strongly dependent on the number of spacetime dimensions. It is trivial to see, for example, that such invariants do not exist in any odd number of dimensions since the number of indices in those is odd and a scalar cannot be formed. In an even number of dimensions, such invariants exist up until $D=8$ because this is the maximum number of free indices that two generalized curvature tensors may have. Therefore, the absence of such invariants in $10D$ heterotic string action should not be a surprise, since they simply do not exist. In $D=4$, the number of parity-odd invariants is four [39]: